Suppose Jonathan breeds flowers and wants to optimize production of offspring with both short stems and white flowers, which are coded for by two genes with the recessive alleles t and p, respectively. In flowers, T codes for tall stems and P codes for purple flowers. Jonathan crosses two heterozygotes that produce 544 offspring. How many of these 544 offspring are predicted to have both short stems and white flowers?
Answer #1

Answer:

TtPp x TtPp ----Parents

Tt x Tt = T_ (tall-3/4) & tt (short-1/4)

Pp x Pp = P_(purple-3/4) & pp (white-1/4)

The proportion of short stems and white flowers = ttpp = ¼*1/4= 1/16

Number of short stems and white flowers plants produced out of 544 = 1/16*544 = 34